# Modeling Multi-Task Model Merging as Adaptive Projective Gradient Descent

### v2

@@ 5.1 Experimental Setup
Implementation details. We perform 400 iterations of learning $\Delta$ with a learning rate of $41e-$ 4. The global magnitude of the merging coefficient $\eta$ is set to 0.07 for vision tasks and 0.15 for NLP tasks. The subspace basis size $k$ is simply defined as the rank of each task vector divided by the number of tasks (i.e., 8). We report Spearman’s $\rho$ for STSB and the standard average accuracy (%) for other tasks. Additional information on the experimental setup for model merging can be found in App. B.
Compared baselines. We categorize the baselines into three main groups: Non-Merging methods, Data-Free methods, and Test-Time Adaptation methods. The non-merging category includes individually fine-tuned models and a traditional multi-task learning approach. The traditional MTL trains the base model on all tasks simultaneously, serving as an upper bound for multi-task model merging. The data-free methods we evaluate include Task Arithmetic (Ilharco et al., 2023), Ties-Merging (Yadav et al., 2024), Consensus Merging (Wang et al., 2024), AWD TA (Xiong et al., 2024), PCB-Merging (Du et al., 2024), and Concrete TA (Tang et al., 2023). Lastly, we include TTA methods such as AdaMerging (Yang et al., 2024c) (layer-wise) and Representation Surgery (Yang et al., 2024b). Further details about these baseline methods are provided in App. C.
@@ 5.2 Main Results
Vision tasks.
Tabs. 1 and 2 present the results for the ViT-B/32 and ViT-L/14 architectures, respectively. Methods like Concrete Merging and Ties-Merging address parameter conflicts by eliminating certain neurons during model merging, outperforming baselines such as TA. AdaMerging and AdaMerging++ automatically learn layer-wise merging coefficients on the test set in an unsupervised manner, also demonstrating strong performance. However, despite these advances, all existing model merging methods still show a noticeable gap compared to individually fine-tuned models. AWD also optimizes $\Delta$ but focuses on increasing orthogonality among task vectors, neglecting the performance gap with individually fine-tuned models. In contrast, our proposed doGe is orthogonal to existing merging methods and can complement them. When applied to Task Arithmetic and AdaMerging, significant performance improvements are observed. For instance, on ViT-B/32, Task Arithmetic’s accuracy improves from 69.1% to 80.7% with doGe. For the test-time adaptation method AdaMerging, accuracy increases from 80.1% to 85.9%. On ViT-L/14, AdaMerging achieves 92.6% accuracy after incorporating doGe, nearly matching the 93.5% achieved by Traditional MTL.
Language tasks.
We extend our approach to language models and LoRA fine-tuned models to evaluate its generalizability (Li et al., 2023). Unlike classification tasks, text-to-text generation requires generating coherent outputs rather than merely projecting hidden representations to logits, introducing additional complexity (Yu et al., 2024). Tabs. 3 and 4 present the results on Flan-T5-base and Flan-T5-large models. Given that pre-trained LLMs already exhibit strong multitasking capabilities, the potential for substantial improvement via specialized methods is inherently limited. Nevertheless, our approach achieves the highest performance, even outperforming TTA methods under data-free conditions. On Flan-T5-large, our data-free method achieves an accuracy of 88.0%, closely approaching the performance of individually fine-tuned models at 89.6%. These results highlight the superior generalization ability of our method across diverse models and tasks.
@@ 5.3 Ablation Studies
Generalization and robustness evaluation.
To further assess the generalization and robustness of our approach, we conduct experiments on unseen tasks and corrupted test sets (i.e., out-of-distribution). Tab. 5 presents generalization results on two unseen tasks. On in-domain tasks, our approach (under data-free conditions) performs comparably to AdaMerging, which leverages the test set for adaptation. Notably, on unseen tasks, where no corresponding task vectors were merged, our method outperforms AdaMerging by an average of 2.3%, demonstrating superior generalization. By contrast, TTA methods rely on the test set, which constrains their ability to generalize. Furthermore, Tab. 8 in Appendix evaluates each method’s robustness on corrupted test sets, designed to simulate real-world scenarios where input data may be noisy or corrupted. The results underline the overall strength and efficacy of our approach, particularly in handling noise and out-of-distribution data.
Effects of each module.
@@
Tab. 6 evaluates the contribution of each module to overall performance. We start with doGe TA and remove one component at a time, reporting the performance for full model merging (ViT-B/32) and for merging PEFT models (T5-base on GLUE). Removing $\Delta$ optimization corresponds to using the task-aware $\lambda$ on TA, underscoring the effectiveness of the data-free objective applied to task vectors, which reduces conflicts between tasks. In cases where the shared subspace is removed, $\Delta$ optimization occurs in the original parameter space. This demonstrates that optimizing within the shared subspace enables the merged model to capture shared knowledge across multiple tasks. When task-aware $\lambda$ is removed, we utilize a uniform merging coefficient of 0.3. Tab. 9 in Appendix further presents the task-wise and layer-wise improvements over TA. Tab. 6 shows that each component is crucial for achieving optimal performance; particularly, $\Delta$ optimization and the shared subspace are most vital, causing notable performance drops of 8.8% and 3.5% in vision tasks, and 2.0% and 0.9% in language tasks, respectively. With all modules included, we achieve the best performance, boosting TA by 5%-11% and demonstrating the complementarity of these components.
@@
Since the effectiveness of our method hinges on the decomposition of the subspace, we explore the impact of the rank ($k$ of $S_{\text{share}}$) on merging performance. Fig. 4 displays the performance with varying rank ratios, alongside the explained standard deviation (i.e., the ratio of preserved singular values $\sigma$ to the total sum of singular values $\Sigma$). Updates performed orthogonally to the subspace direction have shown positive results, with the optimal rank identified between 10%-30%, where the explained standard deviation already exceeds 40%. Preserving a higher rank introduces noise, resulting in a high volume of constraints in the gradient space. Updates along the direction of the shared subspace also slightly outperform those in the original parameter space, due to the allowance for learning personalized subspaces. Tab. 7 reports the specific performance across eight tasks at the same rank. Compared to w/o or updates only along the subspace direction, we observe significant improvements on the DTD dataset but a decrease in performance on the SVHN dataset. This is attributable to DTD’s requirement for rich textural and geometric features, which are well-preserved in the shared subspace. Conversely, SVHN (Street View House Numbers) differs significantly in visual representation from other tasks, making the primary components in the shared subspace less suitable for SVHN. This is further evidenced by the gap from the pre-trained model to individual performance: SVHN shows the lowest pre-trained performance at 31.4%, yet finetuning results peak at 97.5%, indicating a need for task-specific features. In summary, this demonstrates that our method effectively preserves shared knowledge across multiple tasks, achieving optimal overall performance.
@@ 6 Conclusion
The use of large-scale image datasets often involves privacy, labor, and ethical challenges, limiting research opportunities. As a result, the research community is turning towards leveraging pre-trained models. Model merging offers a novel approach to multi-task learning by utilizing the abundant expert models made available by the open-source ethos. With more than a million models accessible on Hugging Face, this strategy leverages the vast resources of the community. This shift enables the creation of multi-task models by directly merging independently trained expert models without needing original training data, presenting a new paradigm.
@@ Appendix A Model Details
For vision tasks, we employ pre-trained models from CLIP (Radford et al., 2021), fine-tuning them using the AdamW optimizer with a weight decay of 0.1 and a learning rate of $1\times 10^{-5}$. The downstream challenges we address include SUN397 (Xiao et al., 2016), Stanford Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2019), SVHN (Netzer et al., 2011), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), and DTD (Cimpoi et al., 2014). We measure the models’ performance using top-1 accuracy as the primary metric.
For NLP tasks, our pre-trained model is Flan-T5 (Chung et al., 2024). We deploy Flan-T5 on eight tasks from the GLUE benchmark (Wang et al., 2019), including CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST2, and STSB. To ensure consistency and reproducibility, we use the same parameter-efficient models following FusionBench (Tang et al., 2024a). The Flan-T5 models, which are encoder-decoder Transformer models, undergo LoRA fine-tuning with hyperparameters $r=16$ and $\alpha=32$ (Hu et al., 2022). We maintain a constant learning rate of $4\times 10^{-5}$ and a uniform batch size of 16 across all tasks, fine-tuning for 2000 steps per task. Adapting to the text-to-text framework, we have restructured the initial inputs accordingly. Performance is evaluated using exact match accuracy for all tasks, except for STSB where we report Spearman’s $\rho$.
@@ Appendix B Implementation Details
The experiments in our study were conducted on a consistent hardware setup, utilizing NVIDIA GTX 4090 GPUs equipped with 24GB of memory. We performed 400 iterations of learning $\Delta$ with a learning rate of $41e-$ 4 using the Adam optimizer. The global magnitude of the merging coefficient $\eta$ is set to 0.07 for vision tasks and 0.15 for NLP tasks. We did not perform a specialized grid search. This setting was chosen because the calculated average $\lambda$ was close to 0.3, which is a beneficial scaling coefficient for the Task Arithmetic method, demonstrating that our approach is not tricky. The subspace basis size $k$ is simply defined as the rank of the task vector divided by the number of tasks (i.e.., 8), with the shared subspace basis size set at the rank divided by 6. For the implementation of our experiments, we employed PyTorch version 2.5 with Python 3.10.
@@ Appendix C Compared Baselines
Ties-Merging (Yadav et al., 2024): Combines steps like trimming, parameter sign determination, and disjoint merging to produce a merged task vector $\boldsymbol{\tau}$. The final model is defined as $\boldsymbol{\theta}=\boldsymbol{\theta_{0}}+\lambda\boldsymbol{\tau}$, where $\lambda$ is tuned using a validation set.
Consensus Merging (Wang et al., 2024): Improves traditional merging methods by removing “selfish” and “catastrophic” weights—parameters beneficial only to specific tasks but detrimental to others.
@@ Appendix D Experiment Results
Robustness.
To evaluate our approach’s robustness to real-world variations, where data characteristics can significantly differ, we conducted extensive ablation studies across diverse data distributions. These studies specifically assessed the model’s performance on out-of-distribution (OOD) data. To simulate real-world conditions, we introduced various types of noise into the test data following the procedure outlined by Yang et al. (2024c). Eight distinct noise types were used—motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast, and JPEG compression—to reflect a wide range of potential distortions encountered in practical applications.
@@
Computational requirements.
As illustrated in Tab. 10, our approach, involves optimizing $\Delta$ in the subspace across 8 vision tasks over 400 iterations. The results show that our approach incurs minimal computational overhead across different model versions and requires only modest GPU memory, facilitated by layer-wise optimization while achieving rapid gradient descent. This demonstrates near-universal scalability on devices equipped with contemporary GPUs.
Model Solving Time GPU Memory
ViT-B/32 121s 729MB
ViT-L/14 311s 2448MB
@@
Fig. 1 illustrates the robustness of our approach when handling a larger number of tasks. Following Wang et al. (2024), we evaluate its performance as more tasks are merged. In addition to the previously used 8 tasks, the 14-task scenario incorporates CIFAR100, STL10, Flowers102, OxfordIIITPet, PCAM, and FER2013. The 20-task scenario further adds six tasks: EMNIST, CIFAR10, Food101, FashionMNIST, RenderedSST2, and KMNIST. Our approach exhibits increasingly significant performance advantages as the number of tasks grows, demonstrating its effectiveness in mitigating negative transfer through gradient descent while preserving task-specific knowledge.

### v3

@@ 5.1 Experimental Setup
Implementation details. We perform 400 iterations of learning $\Delta$ with a learning rate of $41e-$ 4. The global magnitude of the merging coefficient $\eta$ is set to 0.07 for vision tasks and 0.15 for NLP tasks. The subspace basis size $k$ is simply defined as the rank of each task vector divided by the number of tasks (i.e., 8). Following Ties-Merging (Yadav et al., 2023), we retain only the top 30% of parameters with the largest magnitudes. We report Spearman’s $\rho$ for STSB and the standard average accuracy (%) for other tasks. Additional information on the experimental setup for model merging can be found in App. B.
Compared baselines. We categorize the baselines into three main groups: Non-Merging methods, Data-Free methods, and Test-Time Adaptation methods. The non-merging category includes individually fine-tuned models and a traditional multi-task learning approach. The traditional MTL trains the base model on all tasks simultaneously, serving as an upper bound for multi-task model merging. The data-free methods we evaluate include Task Arithmetic (Ilharco et al., 2023), Ties-Merging (Yadav et al., 2023), Consensus Merging (Wang et al., 2024b), AWD TA (Xiong et al., 2024), PCB-Merging (Du et al., 2024), and Concrete TA (Tang et al., 2023). Lastly, we include TTA methods such as AdaMerging (Yang et al., 2024c) (layer-wise) and Representation Surgery (Yang et al., 2024b). Further details about these baseline methods are provided in App. C.
@@ 5.2 Main Results
Vision tasks. Tabs. 1 and 2 present the results for the ViT-B/32 and ViT-L/14 architectures, respectively. Methods like Concrete Merging and Ties-Merging address parameter conflicts by eliminating certain neurons during model merging, outperforming baselines such as TA. AdaMerging and AdaMerging++ automatically learn layer-wise merging coefficients on the test set in an unsupervised manner, also demonstrating strong performance. However, despite these advances, all existing model merging methods still show a noticeable gap compared to individually fine-tuned models. AWD also optimizes $\Delta$ but focuses on increasing orthogonality among task vectors, neglecting the performance gap with individually fine-tuned models. In contrast, our proposed doGe is orthogonal to existing merging methods and can complement them. When applied to Task Arithmetic and AdaMerging, significant performance improvements are observed. For instance, on ViT-B/32, Task Arithmetic’s accuracy improves from 69.1% to 80.7% with doGe. For the test-time adaptation method AdaMerging, accuracy increases from 80.1% to 85.9%. On ViT-L/14, AdaMerging achieves 92.6% accuracy after incorporating doGe, nearly matching the 93.5% achieved by Traditional MTL.
Language tasks. We extend our approach to language models and LoRA fine-tuned models to evaluate its generalizability (Li et al., 2023). Unlike classification tasks, text-to-text generation requires generating coherent outputs rather than merely projecting hidden representations to logits, introducing additional complexity (Li et al., 2025b). Tabs. 3 and 4 present the results on Flan-T5-base and Flan-T5-large models. Given that pre-trained LLMs already exhibit strong multitasking capabilities, the potential for substantial improvement via specialized methods is inherently limited. Nevertheless, our approach achieves the highest performance, even outperforming TTA methods under data-free conditions. On Flan-T5-large, our data-free method achieves an accuracy of 88.0%, closely approaching the performance of individually fine-tuned models at 89.6%. These results highlight the superior generalization ability of our method across diverse models and tasks.
@@ 5.3 Ablation Studies
Generalization and robustness evaluation. To further assess the generalization and robustness of our approach, we conduct experiments on unseen tasks and corrupted test sets (i.e., out-of-distribution). Tab. 5 presents generalization results on two unseen tasks. On in-domain tasks, our approach (under data-free conditions) performs comparably to AdaMerging, which leverages the test set for adaptation. Notably, on unseen tasks, where no corresponding task vectors were merged, our method outperforms AdaMerging by an average of 2.3%, demonstrating superior generalization. By contrast, TTA methods rely on the test set, which constrains their ability to generalize. Furthermore, Tab. 11 in Appendix evaluates each method’s robustness on corrupted test sets, designed to simulate real-world scenarios where input data may be noisy or corrupted. The results underline our approach’s overall strength and efficacy, particularly in handling noise and out-of-distribution data.
@@
Effects of each module. Tab. 6 evaluates the contribution of each module to overall performance. We start with doGe TA and remove one component at a time, reporting the performance for full model merging (ViT-B/32) and for merging PEFT models (T5-base on GLUE). Removing $\Delta$ optimization corresponds to using the task-aware $\lambda$ on TA, underscoring the effectiveness of the data-free objective applied to task vectors, which reduces conflicts between tasks. In cases where the shared subspace is removed, $\Delta$ optimization occurs in the original parameter space. This demonstrates that optimizing within the shared subspace enables the merged model to capture shared knowledge across multiple tasks. When task-aware $\lambda$ is removed, we utilize a uniform merging coefficient of 0.3. Tab. 13 in Appendix further presents the task-wise and layer-wise improvements over TA. Tab. 6 shows that each component is crucial for achieving optimal performance; particularly, $\Delta$ optimization and the shared subspace are most vital, causing notable performance drops of 8.8% and 3.5% in vision tasks, and 2.0% and 0.9% in language tasks, respectively. With all modules included, we achieve the best performance, boosting TA by 5%-11% and demonstrating the complementarity of these components.
@@
Effects of the subspace. Since the effectiveness of our method hinges on the decomposition of the subspace, we explore the impact of the rank ($k$ of $S_{\text{share}}$) on merging performance. Fig. 4 displays the performance with varying rank ratios, alongside the explained standard deviation (i.e., the ratio of preserved singular values $\sigma$ to the total sum of singular values $\Sigma$). Updates performed orthogonally to the subspace direction have shown positive results, with the optimal rank identified between 10%-30%, where the explained standard deviation already exceeds 40%. Preserving a higher rank introduces noise, resulting in a high volume of constraints in the gradient space. Updates along the direction of the shared subspace also slightly outperform those in the original parameter space, due to the allowance for learning personalized subspaces. Tab. 7 reports the specific performance across eight tasks at the same rank. Compared to w/o or updates only along the subspace direction, we observe significant improvements on the DTD dataset but decreased performance on the SVHN dataset. This is attributable to DTD’s requirement for rich textural and geometric features, which are well-preserved in the shared subspace. Conversely, SVHN (Street View House Numbers) differs significantly in visual representation from other tasks, making the primary components in the shared subspace less suitable for SVHN. This is further evidenced by the gap from the pre-trained model to individual performance: SVHN shows the lowest pre-trained performance at 31.4%, yet finetuning results peak at 97.5%, indicating a need for task-specific features. In summary, this demonstrates that our method effectively preserves shared knowledge across multiple tasks, achieving optimal overall performance.
Hyperparameter sensitivity. Additional sensitivity analysis for the global scaling factor $\eta$ is provided in Tab. 8. Evaluations across $\eta$ values from 0.01 to 0.09 show that performance remains stable, even reaching higher values at certain points. (We did not conduct an exhaustive grid search; this range was chosen because the computed $\eta$ was close to 0.03.) This consistent performance across different $\eta$ values demonstrates the robustness of our approach and highlights the practicality of task-aware coefficients.
Table 8: Sensitivity analysis for the global scaling factor $\eta$.
$\eta$ 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09
ViT-B/32 79.5 80.3 80.6 80.9 81.0 80.8 80.7 80.2 79.8
Model Solving Time GPU Memory
ViT-B/32 121s 729MB
ViT-L/14 311s 2448MB
Computational requirements. As illustrated in Tab. 9, our approach involves optimizing $\Delta$ within the subspace across 8 vision tasks over 400 iterations. The results demonstrate that our method incurs minimal computational overhead across different model variants and requires only moderate GPU memory. This efficiency is achieved through layer-wise optimization and fast convergence via gradient descent. Notably, the SVD operation is performed only once at the beginning, with a computational complexity of $O(\min(mn^{2},m^{2}n))$. These findings highlight the near-universal scalability of our method on devices equipped with modern GPUs.
Generative language tasks. We further extend our method to LLMs and conduct experiments following standard settings (Yu et al., 2024). The merging process is completed in just 58 minutes on a single A100 GPU. We report normalized scores relative to the performance of individual models when merging WizardLM-13B (Instruction-Following), WizardMath-13B (Math), and llama-2-13b-code-alpaca (Code). As shown in Tab. 10, our method achieves the highest average performance across tasks, demonstrating its effectiveness and scalability in generative language tasks.
Table 10: Normalized scores are computed relative to individual models when merging WizardLM-13B (Instruction-Following), WizardMath-13B (Math), and LLaMA-2-13B-code-alpaca (Code).
Method AlpacaEval GSM8K MATH HumanEval MBPP Avg.
Individual 100.0 100.0 100.0 100.0 100.0 100.0
TA 102.7 91.0 70.5 50.0 87.7 80.4
TIES 98.1 97.4 68.1 60.0 89.4 82.6
TA + DARE 103.1 88.0 72.5 63.3 92.9 84.0
TIES + DARE 107.9 90.3 65.6 80.0 92.4 87.2
Ours 107.5 105.0 94.4 56.7 86.5 90.0
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012, KJZD20230923114916032, GJHZ20240218113604008), Beijing Key Lab of Networked Multimedia, the Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO. JCYJ20241202124430041), National Natural Science Foundation of China (No. 62025604).
@@ Acknowledgments
The use of large-scale image datasets often involves privacy, labor, and ethical challenges, limiting research opportunities. As a result, the research community is turning towards leveraging pre-trained models. Model merging offers a novel approach to multi-task learning by utilizing the abundant expert models made available by the open-source ethos. With more than a million models accessible on Hugging Face, this strategy leverages the community’s vast resources. This shift enables the creation of multi-task models by directly merging independently trained expert models without needing original training data, presenting a new paradigm.
@@ Appendix A Model Details
For vision tasks, we employ pre-trained models from CLIP (Radford et al., 2021), fine-tuning them using the AdamW optimizer with a weight decay of 0.1 and a learning rate of $1\times 10^{-5}$. The downstream tasks encompass a variety of challenges. SUN397 (Xiao et al., 2016) is a large-scale scene recognition dataset comprising over 100,000 images across 397 indoor and outdoor scene categories. Stanford Cars (Krause et al., 2013) contains 16,185 images of 196 car models and is commonly used for fine-grained image classification. RESISC45 (Cheng et al., 2017) consists of 31,500 remote sensing images evenly distributed over 45 scene categories, supporting research in aerial scene classification. EuroSAT (Helber et al., 2019) is based on Sentinel-2 satellite images and includes 27,000 samples covering 10 land use and land cover classes. SVHN (Netzer et al., 2011) is a real-world digit recognition dataset with over 600,000 images of house numbers captured from Google Street View. GTSRB (Stallkamp et al., 2011) comprises more than 50,000 images of 43 traffic sign categories, serving as a benchmark for traffic sign recognition tasks. MNIST (LeCun, 1998) is a well-known dataset for handwritten digit classification, featuring 70,000 grayscale images of digits from 0 to 9. DTD (Cimpoi et al., 2014) is a texture dataset with 5,640 images organized into 47 human-describable categories, designed for studying texture perception and classification. We measure the models’ performance using top-1 accuracy as the primary metric (Horoi et al., 2024; Stoica et al., 2024; Wei et al., 2025).
For NLP tasks, our pre-trained model is Flan-T5 (Wang et al., 2024a). We deploy Flan-T5 on eight tasks from the GLUE benchmark (Wang et al., 2019), including CoLA, MNLI, MRPC, QNLI, QQP, RTE, SST2, and STSB. To ensure consistency and reproducibility, we use the same parameter-efficient models following FusionBench (Tang et al., 2024a). The Flan-T5 models, which are encoder-decoder Transformer models, undergo LoRA fine-tuning with hyperparameters $r=16$ and $\alpha=32$ (Hu et al., 2022). We maintain a constant learning rate of $4\times 10^{-5}$ and a uniform batch size of 16 across all tasks, fine-tuning for 2000 steps per task. Adapting to the text-to-text framework, we have restructured the initial inputs accordingly. Performance is evaluated using exact match accuracy for all tasks, except for STSB where we report Spearman’s $\rho$.
@@ Appendix B Implementation Details
The experiments in our study were conducted on a consistent hardware setup, utilizing NVIDIA GTX 4090 GPUs equipped with 24GB of memory. We performed 400 iterations of learning $\Delta$ with a learning rate of $41e-$ 4 using the Adam optimizer. The global magnitude of the merging coefficient $\eta$ is set to 0.07 for vision tasks and 0.15 for NLP tasks. We did not perform a specialized grid search. This setting was chosen because the calculated average $\lambda$ was close to 0.3, which is a beneficial scaling coefficient for the Task Arithmetic method, demonstrating that our approach is not tricky. The subspace basis size $k$ is simply defined as the rank of the task vector divided by the number of tasks (i.e.., 8), with the shared subspace basis size set at the rank divided by 6. Following Ties-Merging (Yadav et al., 2023), we retain only the top 30% of parameters with the largest magnitudes. We only apply our method to the linear layer in the model. For the implementation of our experiments, we employed PyTorch version 2.5 with Python 3.10.
@@ Appendix C Compared Baselines
Ties-Merging (Yadav et al., 2023): Combines steps like trimming, parameter sign determination, and disjoint merging to produce a merged task vector $\boldsymbol{\tau}$. The final model is defined as $\boldsymbol{\theta}=\boldsymbol{\theta_{0}}+\lambda\boldsymbol{\tau}$, where $\lambda$ is tuned using a validation set.
Consensus Merging (Wang et al., 2024b): Improves traditional merging methods by removing “selfish” and “catastrophic” weights—parameters beneficial only to specific tasks but detrimental to others.
@@ Appendix D Experiment Results
Robustness.
To evaluate our approach’s robustness to real-world variations, where data characteristics can significantly differ, we conducted extensive ablation studies across diverse data distributions. These studies specifically assessed the model’s performance on out-of-distribution (OOD) data (Zhang et al., 2024a, b; Dong et al., 2023a, b; Zhang et al., 2025). To simulate real-world conditions, we introduced various types of noise into the test data following the procedure outlined by Yang et al. (2024c). Eight distinct noise types were used—motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast, and JPEG compression—to reflect a wide range of potential distortions encountered in practical applications.
We conduct experiments evaluating generalization on three unseen tasks when merging five other tasks. The results in Tab. 12 reveal that SUN397, DTD, and Cars datasets pose challenges for ViT models, while MNIST/EuroSAT show limited generalization to these complex tasks. Despite this, our method consistently outperformed other model merging approaches by a significant margin.
Table 12: Generalization results on three unseen tasks when merging ViT-B/32 models on five tasks.
Method Seen Tasks Unseen Tasks
RESISC45 SVHN GTSRB MNIST EuroSAT Avg. SUN397 Cars DTD Avg.
Pre-trained 60.6 23.5 30.4 47.6 45.6 41.5 63.2 59.9 43.9 55.6
Task Arithmetic 52.8 83.9 71.1 97.7 61.9 73.5 27.9 25.0 26.4 26.4
Ties-Merging 74.6 89.1 81.8 97.7 73.7 83.4 57.5 51.9 38.7 49.4
AdaMerging 73.5 76.0 81.5 97.4 69.4 79.6 42.3 37.8 32.0 37.4
doGe TA (Ours) 82.6 89.4 89.0 98.6 92.3 90.4 58.7 54.3 41.4 51.5
@@
Tab. 14 illustrates the robustness of our approach when handling a larger number of tasks. Following Wang et al. (2024b), we evaluate its performance as more tasks are merged. In addition to the previously used 8 tasks, the 14-task scenario incorporates CIFAR100, STL10, Flowers102, OxfordIIITPet, PCAM, and FER2013. The 20-task scenario further adds six tasks: EMNIST, CIFAR10, Food101, FashionMNIST, RenderedSST2, and KMNIST. Our approach exhibits increasingly significant performance advantages as the number of tasks grows, demonstrating its effectiveness in mitigating negative transfer through gradient descent while preserving task-specific knowledge.
@@
Comparisons with dynamic merging.
As shown in Tab. 15, merging multiple models into a single model presents notable challenges. doGe is a static, plug-and-play merging method (similar to Task Arithmetic) that maintains the standard model size and supports parallelized inference. In contrast, dynamic merging approaches (Tang et al., 2024b; Huang et al., 2024; Lu et al., 2024) offer greater flexibility by dynamically selecting task-specific modules but typically require additional storage and encounter scalability considerations during inference. These methods often rely on either dynamic I/O loading of modules or maintaining all components in GPU memory. For instance, some methods train routing networks using validation data to guide module selection.
Table 15: Distinction based on parameters, data requirements, and computational costs.
Method Parameters Router Data Parallel Performance
Task Arithmetic 1×1$\times$1 × - - static 69.1
AdaMerging 1×1$\times$1 × - unlabeled test dataset static 80.1
doGe TA 1×1$\times$1 × - - static 81.0 ($\uparrow$11.6)
doGe AM 1×1$\times$1 × - unlabeled test dataset static 85.9 ($\uparrow$5.8)
Representation Surgery >1×>1$\times$> 1 × - unlabeled test dataset static 80.9
EMR merging 4×4$\times$4 × perfect router - dynamic 88.7
Twin merging 2.25×2.25$\times$2.25 × trained router labeled validation dataset dynamic 86.1
WEMoE 5×5$\times$5 × trained router unlabeled test dataset dynamic 89.4
Traditional MTL 1×1$\times$1 × - - - 88.9
Multiple Models 8×8$\times$8 × - - - 90.8
Potential limitations
A potential limitation is the lack of consideration for heterogeneous model merging, which requires transformation when task vectors have inconsistent shapes or layer numbers.